KingPizza:

(Calculus) What would be the normal vector to 4x+6y+5z=2?

jayfafr:

4x+6y+5z=2 is ⟨4,6,5⟩

jayfafr:

@kingpizza

wio:

The normal vector to a plane is a vector that is perpendicular to the plane. In \(\mathbb R^3\) it is easy to find the normal vector to a plane from an equation, because it is just the coefficients of the variables, or in this case \(\langle 4,6,5\rangle\). But why is this the case? To form the equation of a plane in \(\mathbb R^3\) we use a normal vector \(\vec n\) and some point \(P_0\) which lies on the plane. We can test if any arbitrary point \(P = (x,y,z) \) lies on the plane, because the vector \(\vec r = P-P_0\) must be parallel to the plane and orthogonal to \(\vec n\). So \(\vec n \cdot \vec r = 0\). Suppose be didn't know the equation \(4x+6y+5z=2\), we just knew \(\vec n = \langle 4,6,5\rangle\) and \(P_0 = (-1,1,0)\). Then we can use the vector equation to find the plane equation: \[ \begin{array}{rcl} \vec n \cdot \big( P - P_0 \big) &=& 0 \\ \langle 4,6,5\rangle \cdot \big( (x,y,z) - (-1,1,0) \big) &=& 0 \\ \langle 4,6,5\rangle \cdot \langle x+1,y-1,z\rangle&=& 0 \\ 4(x+1) + 6(y-1) + 5z &=& 0 \\ 4x + 6y + 5z -2 &=& 0 \\ 4x + 6y + 5z &=& 2 \\ \end{array} \] All equations of planes in \(\mathbb R^3\) have the form \(ax+by+cz = d\).

pup:

i love math but ill try this ima say its 2